For the next 6 questions suppose the following data Cash & M/S Receivables Inventories Fixed Assets Total Assets Balance Sheet 30,000 Accruals 70,000 Accounts Payable 120,000 Notes Payable 480,000L-TLiabilities 700,000 Common Equity 20,000 60,000 80,000 140,000 400 700 Total Annual Sales $1,400,000 Net Income$63,000 Number of Shares Outstanding 30 000 ost of Goods Sold $960,000 What is the current ratio of the firm? 1.225 O 1.375 O 1.875 O 1.905 O 2.145
media%2F1c3%2F1c3dcae1-16f6-4f31-8a60-3f
media%2Fa0f%2Fa0fdb90d-d6ca-4158-b5b7-11
media%2F9a1%2F9a11e819-47f5-42a4-9baf-f1

media%2F372%2F372960e2-17f0-4ba1-b2f7-49
media%2Fbec%2Fbec31647-edbe-4f09-b77b-6f
Answer #1

1. Current Ratio = Current Assets / Current Liabilities

= (Cash + Inventories + Receivables) /( Accounts Payable + Notes payable + Accruals)

= 220,000 / 160,000 = 1.375

2. Debt ratio = Total Debt / Total Assets = 300,000 / 700,000 = 0.43 (rounded off)

3. Asset turnover ratio = Sales / Total assets = 1,400,000 /700,000 = 2

4. Days sales outstanding = [ Accounts Receivable / (Total Credit Sales / No of days)] = [ 70,000/ (1,400,000/360)] = 18 days

5. Profit Margin = (Net Income / Sales) * 100 =( 63,000 / 1,400,000)*100 = 4.5%

6. Return on equity = (Net Income / Common Equity) * 100 = (63,000 / 400,000)*100 = 15.75%